# Quantum Speedups in Regret Analysis of Infinite Horizon Average-Reward Markov Decision Processes

### v3

@@ 1 Introduction
Quantum Machine Learning (QML) has garnered remarkable interest in contemporary research, predominantly attributed to the pronounced speedups achievable with quantum computers as opposed to their classical analogs (?, ?). The fundamental edge of quantum computing arises from the unique nature of its fundamental computing element, termed a qubit, which can exist simultaneously in states of 0 and 1, unlike classical bits that are restricted to either 0 or 1. This inherent distinction underpins the exponential advancements that quantum computers bring to specific computational tasks, surpassing the capabilities of classical computers.
@@
In the realm of Reinforcement Learning (RL), an agent embarks on the task of finding an efficient policy for Markov Decision Process (MDP) environment through repetitive interactions (?). RL has found notable success in diverse applications, including but not limited to autonomous driving, ridesharing platforms, online ad recommendation systems, and proficient gameplay agents (?, ?, ?, ?). Most of these setups require decision making over infinite horizon, with an objective of average rewards. This setup has been studied in classical reinforcement learning (?, ?, ?), where $\tilde{\mathcal{O}}(\sqrt{T})$ regret across $T$ rounds has been found meeting the theoretical lower bound (?). This paper embarks on an inquiry into the potential utilization of quantum statistical estimation techniques to augment the theoretical convergence speeds of tabular RL algorithms within the context of infinite horizon learning settings.
@@
A diverse array of quantum statistical estimation primitives has emerged, showcasing significant enhancements in convergence speeds and gaining traction within Quantum Machine Learning (QML) frameworks (?, ?, ?). Notably, (?) introduces a quantum mean estimation algorithm that yields a quadratic acceleration in sample complexity when contrasted with classical mean estimation techniques. In this work, we emphasize the adoption of this particular mean estimation method as a pivotal component of our methodology. It strategically refines the signals garnered by the RL agent during its interaction with the enigmatic quantum-classical hybrid MDP environment.
@@
It is pertinent to highlight a fundamental element in the analysis of conventional Reinforcement Learning (RL), which involves the utilization of martingale convergence theorems. These theorems play a crucial role in delineating the inherent stochastic process governing the evolution of states within the Markov Decision Process (MDP). Conversely, this essential aspect lacks parallelism within the quantum setting, where comparable martingale convergence results remain absent. To address this disparity, we introduce an innovative approach to regret bound analysis for quantum RL. Remarkably, our methodology circumvents the reliance on martingale concentration bounds, a staple in classical RL analysis, thus navigating uncharted territories in the quantum realm.
@@
    Our paper introduces a pioneering contribution in the form of Quantum-UCRL (Q-UCRL), an innovative model-based, infinite horizon, optimism-driven quantum Reinforcement Learning (QRL) algorithm. Drawing inspiration from classical predecessors like the UCRL and UC-CURL algorithms (?, ?, ?), Q-UCRL is designed to seamlessly weave together an agent’s optimistic policy acquisition and an adept quantum mean estimator.
@@
    Employing meticulous theoretical analysis, we show that the Q-UCRL algorithm achieves an exponential improvement in the regret analysis. Specifically, it attains a regret bound of $\tilde{\mathcal{O}}(1)$, breaking through the theoretical classical lower bound of $\Omega(\sqrt{T})$ across $T$ online rounds. The analysis is based on the novel quantum Bellman error based analysis (introduced in classical RL in (?) for optimism based algorithm), where the difference between the performance of a policy on two different MDPs is bounded by long-term averaged Bellman error and the quantum mean estimation is used for improved Bellman error. Further, the analysis avoids dependence on martingale concentration bounds and incorporates the phenomenon of state collapse following measurements, a crucial aspect in quantum computing analysis.
@@
2 Related Work
Infinite Horizon Reinforcement Learning: Regret analysis of infinite horizon RL in the classical setting has been widely studied with average reward objective. A prominent principle that underpins algorithms tailored for this scenario is the concept of “optimism in the face of uncertainty” (OFU). In this approach, the RL agent nurtures optimistic estimations of value functions and, during online iterations, selects policies aligned with the highest value estimates (?, ?, ?). Additionally, it’s noteworthy to acknowledge that several methodologies are rooted in the realm of posterior sampling, where the RL agent samples an MDP from a Bayesian Distribution and subsequently enacts the optimal policy (?, ?, ?, ?). In our study, we embrace the OFU-based algorithmic framework introduced in (?), and we extend its scope to an augmented landscape where the RL agent gains access to supplementary quantum information. Furthermore, we render a mathematical characterization of regret, revealing a $\tilde{\mathcal{O}}(1)$ bound, which in turn underscores the merits of astutely processing the quantum signals within our framework.
Quantum Mean Estimation: The realm of mean estimation revolves around the identification of the average value of samples stemming from an unspecified distribution. Of paramount importance is the revelation that quantum mean estimators yield a quadratic enhancement when juxtaposed against their classical counterparts (?, ?). The key reason for this improvement is based on the quantum amplitude amplification, which allows for suppressing certain quantum states w.r.t. the states that are desired to be extracted (?). In Appendix B, we present a discussion around Quantum Amplitude Estimation and its applications in QML.
@@ 2 Related Work
Definition 1 (Random Variable, Definition 2.2 of (?)).
@@
The unitiary $\mathcal{U}_{\mathbb{P}}$ provides the ability to query samples of the random variable in superpositioned states, which is the essence for speedups in quantum algorithms. To perform quantum mean estimation for values of random variables (?, ?, ?), an additional quantum evaluation oracle would be needed.
@@
Next, we present a key quantum mean estimation result that will be carefully employed in our algorithmic framework to utilize the quantum superpositioned states collected by the RL agent. One of the crucial aspects of Lemma 1 is that quantum mean estimation converges at the rate $\mathcal{O}(\frac{1}{n})$ as opposed to classical benchmark convergence rate $\mathcal{O}(\frac{1}{\sqrt{n}})$ for $n$ number of samples, therefore estimation efficiency quadratically.
Lemma 1 (Quantum multivariate bounded estimator, Theorem 3.3 of (?)).
Let $X$ be a d-dimensional bounded random variable such that $||X||_{2}\leq 1$. Given three reals $L_{2}\in(0,1]$, $\delta\in(0,1)$ and $n\geq 1$ such that $\mathbb{E}[||X||_{2}]\leq L_{2}$, the multivariate bounded estimator $\texttt{QBounded}_{d}(X,L_{2},n,\delta)$ obtained by Algorithm 1 outputs a mean estimate $\hat{\mu}$ of $\mu=\mathbb{E}[X]$ such that,
@@
The algorithm performs $\mathcal{O}\Big{(}n\log^{1/2}(n\sqrt{d})\Big{)}$ quantum experiments.
Algorithm 1 $\texttt{QBounded}_{d}(X,L_{2},n,\delta)$ (Algorithm 1 in (?))
@@
8:     Compute the state $|\psi\rangle:=\widetilde{P}_{X,L_{2},m,\alpha,\epsilon}|G\rangle$ in $H_{G}\otimes H_{aux}$, where $\widetilde{P}_{X,L_{2},m,\alpha,\epsilon}$ is the directional mean oracle constructed by the quantum evaluation oracle in Proposition 3.2 (?) with $\epsilon=1/25$.
@@
Quantum Reinforcement Learning: Within the realm of QRL, a prominent strand of previous research showcases exponential enhancements in regret through amplitude amplification-based methodologies applied to the Quantum Multi-Armed Bandits (Q-MAB) problem (?, ?). Nevertheless, the theoretical underpinnings of the aforementioned Q-MAB approaches do not seamlessly extend to the QRL context since there is no state evolution in bandits.
A recent surge of research interest has been directed towards Quantum Reinforcement Learning (QRL) (?, ?, ?). It is noteworthy to emphasize that the aforementioned studies do not provide an exhaustive mathematical characterization of regret. (?) delves into QRL within the infinite horizon Markov Decision Process (MDP) learning framework, yet their focus rests on a discounted reward setup, where advancements in sample complexity are demonstrated through the lens of generative model assumptions. Further, finite horizon episodic MDP setup has been considered in (?). Ours is the first paper to study infinite horizon MDP with average reward objective.
3 Problem Formulation
@@
Quantum Computing Basics: In this section, we give a brief introduction to the concepts that are most related to our work based on (?) and (?). Consider an $m$ dimensional Hilbert space $\mathbb{C}^{m}$, a quantum state $|x\rangle=(x_{1},...,x_{m})^{T}$ can be seen as a vector inside the Hilbert space with $\sum_{i}|x_{i}|^{2}=1$. Furthermore, for a finite set with $m$ elements $\xi=\{\xi_{1},...,\xi_{m}\}$, we can assign each $\xi_{i}\in\xi$ with a member of an orthonormal basis of the Hilbert space $\mathbb{C}^{m}$ by
@@
Quantum transition oracle: We utilize the concepts and notations of quantum computing in (?, ?, ?, ?) to construct the quantum sampling oracle of RL environments and capture agent’s interaction with the unknown MDP environment. We now formulate the equivalent quantum-accessible RL environments for our classical MDP $\mathcal{M}$. For an agent at step $t$ in state $s_{t}$ and with action $a_{t}$, we construct the quantum sampling oracles for variables of the next state $P(\cdot|s_{t},a_{t})$. Specifically, suppose there are two Hilbert spaces $\mathcal{\bar{S}}=\mathbb{C}^{|\mathcal{S}|}$ and $\mathcal{\bar{A}}=\mathbb{C}^{|\mathcal{A}|}$ containing the superpositions of the classical states and actions. We represent the computational bases for $\mathcal{\bar{S}}$ and $\mathcal{\bar{A}}$ as $\{|s\rangle\}_{s\in\mathcal{S}}$ and $\{|a\rangle\}_{a\in\mathcal{A}}$. We assume that we can implement the quantum sampling oracle $\mathcal{U}_{P}$ of the MDP’s transitions as follows:
@@
    $\mathcal{U}_{P}:|s_{t}\rangle\otimes|a_{t}\rangle\otimes|0\rangle\rightarrow|\bm{\psi}\rangle^{t}$
    Where $|\bm{\psi}\rangle^{t}=|s_{t}\rangle\otimes|a_{t}\rangle\otimes\sum_{s^{\prime}\in\mathcal{S}}\sqrt{P(s^{\prime}|s_{t},a_{t})}|s^{\prime}\rangle$.
@@
As described in Fig 1, at step $t$, the agent plays an action $a_{t}$ based on current state $s_{t}$ according to its policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$. Consequently, the unknown MDP Q-environment shares the next state information and reward i.e., $\{s_{t+1},r_{t}\}$. Additionally, the quantum transition oracle first encodes $s_{t}$ and $a_{t}$ into the corresponding basis of the quantum Hilbert spaces, producing basis $|s_{t}\rangle$ and $|a_{t}\rangle$, then it encodes the superpositioned quantum state $|\bm{\psi}\rangle^{t}$ from the quantum transition oracle $\mathcal{U}_{P}$. $|\bm{\psi}\rangle^{t}$ would be used for the Quantum Mean Estimator QBounded outlined in Lemma 1 which in turn improves transition probability estimation and leads to exponentially better regret. We note that a single copy of $|\bm{\psi}\rangle^{t}$ is obtained at each $(s_{t},a_{t})$. Given that a measurement collapses the state, we can only perform one measurement on it. Thus, the quantum mean estimator will only perform measurements at the end of epochs, as will be described in the algorithm.
RL regret formulation: Consider the RL agent following some policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$ on the unknown MDP Q-environment $\mathcal{M}$. Then, as a result of playing policy $\pi$, the agent’s long term average reward is defined next according to (?).
@@
Then, according to (?), $\Gamma_{\pi}^{P}$ can be represented as the following:
@@
wherein, $V_{\pi}^{P}(\cdot;\cdot)$ is the discounted cumulative average reward by following policy $\pi$, $\rho_{\pi}^{P}$ is the steady-state occupancy measure (i.e., $\rho_{\pi}^{P}(s,a)\in[0,1],~{}\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\rho_{\pi}^{P}(s,a)=1$), $\bar{r}(s,a)$ is the steady-state long-term average reward of pair $(s,a)$. In the following, we introduce key assumptions crucial from the perspective of our algorithmic framework and its analysis. In this context, we introduce the notations $\{P_{\pi,s}^{t},T_{\pi,~{}s\rightarrow s^{\prime}}\}$ to denote the $t$ time step probability distribution obtained by playing policy $\pi$ on the unknown MDP $\mathcal{M}$ starting from some arbitrary state $s\in\mathcal{S}$; and, the actual time steps to reach $s^{\prime}$ from $s$ by playing policy $\pi$ respectively. With these notations, we introduce our first assumption pertaining to ergodicity of MDP $\mathcal{M}$.
@@
We emphasize that Assumption 2 is a commonly used assumption in RL algorithm development owing to the fact that in most setups rewards are constructed according to the underlying problem and is known beforehand (?, ?). Next, we define the cumulative regret accumulated by the agent in expectation across $T$ time steps as follows:
@@ 4 Algorithmic Framework
However, in the absence of knowledge pertaining to true $P$, we propose to solve an optimistic version of Eq. (11) that will update RL agent’s policy in epochs.
Remark 1.
Algorithm 2 runs in epochs wherein the agent’s action policy $\pi$ is updated only at the beginning of a new epoch, and the policy is held constant throughout the epoch.
Specifically, at the start of an epoch $e$, the following optimization problem will be solved:
@@
Observe that Eq. (15) increases the search space for policy in a carefully computed neighborhood of certain transition probability estimates $\hat{P}$ via Eq. (19) at the start of epoch $e$. Also in Eq. (19), recall that $S=|\mathcal{S}|$ is the number of states, $A=|\mathcal{A}|$ is the number of actions, $N_{e}(s,a)$ is the number of visitations of state-action pair $(s,a)$ upto the end of epoch $e$.
Note that, solving Eq. (15) is equivalent to obtaining an optimistic policy as it extends the search space of original problem Eq. (11). Furthermore, the computation of $\hat{P}_{e}$ is key to our framework design. Since performing quantum mean estimation to quantum samples would require measuring and results in quantum samples collapsing, each quantum sample can only be used once. We emphasize that the quantum mean estimator will only run at the end of the epoch, performing measurements to update the transition probability for the next epoch.
Details on Quantum Mean Estimator: At the end of the epoch, we use all $(s_{i},a_{i})$ observed in the samples, and perform quantum mean estimation on all the samples $|\bm{\psi}\rangle^{t}$ obtained for that $(s_{i},a_{i})$. More precisely, $\tilde{P}_{e}(\cdot|s_{i},a_{i})$ is the quantum mean estimation for the random variable $P_{e}(\cdot|s_{i},a_{i})$ performed by QBounded (Algorithm 1 in (?)) using the quantum samples $\{|\bm{\psi}\rangle\}^{e}_{i}$ observed in the previous epoch with the total number being $\nu_{e-1}(s_{i},a_{i})$:
${\{|\bm{\psi}\rangle\}^{e}_{i}=\{|\bm{\psi}\rangle^{t}\,|\;t_{e-1}^{start}\leq t\leq t_{e}^{start}-1,(s_{t},a_{t})=(s_{i},a_{i})\}}$ (20)
@@
for some $c\in\mathbb{R}$ such that the term $cn\log^{1/2}(n\sqrt{S})$ is greater than the number of quantum experiments needed for performing Eq. (21), which is $\mathcal{O}\Big{(}n\log^{1/2}(n\sqrt{S})\Big{)}$. In addition, since $\mathbb{E}[||P_{e}(\cdot|s_{i},a_{i})||_{2}]\leq 1$, we set $L_{2}=1.$ Different from the works in classical approaches based on (?), our algorithm utilizes each quantum sample only once for estimating all $\hat{P}_{e}$ throughout the entire process. This strategy fits with our quantum framework since quantum samples would collapse once being measured. In Appendix A, we show that Eq. (15) is a convex optimization problem, therefore standard optimization solvers can be employed to solve it. Finally, the agent’s policy at the start of epoch $e$ is given by:
@@
Next, we formally present our method Q-UCRL in Algorithm 2.
@@
11:     Observe $s_{t+1},r(s_{t},a_{t})$ and quantum samples $|\bm{\psi}\rangle^{t}$.

### v4

# Washington, DC 20004 USA

@@ Introduction
Quantum Machine Learning (QML) has garnered remarkable interest in contemporary research, predominantly attributed to the pronounced speedups achievable with quantum computers as opposed to their classical analogs (biamonte2017quantum; bouland2023quantum). The fundamental edge of quantum computing arises from the unique nature of its fundamental computing element, termed a qubit, which can exist simultaneously in states of 0 and 1, unlike classical bits that are restricted to either 0 or 1. This inherent distinction underpins the exponential advancements that quantum computers bring to specific computational tasks, surpassing the capabilities of classical computers.
@@
In the realm of Reinforcement Learning (RL), an agent embarks on the task of finding an efficient policy for Markov Decision Process (MDP) environment through repetitive interactions (sutton2018reinforcement). RL has found notable success in diverse applications, including but not limited to autonomous driving, ridesharing platforms, online ad recommendation systems, and proficient gameplay agents (silver2017mastering; al2019deeppool; chen2021deepfreight; bonjour2022decision). Most of these setups require decision making over infinite horizon, with an objective of average rewards. This setup has been studied in classical reinforcement learning (auer2008near; agarwal2021concave; fruit2018efficient), where $\tilde{\mathcal{O}}(\sqrt{T})$ regret across $T$ rounds has been found meeting the theoretical lower bound (jaksch2010near). This paper embarks on an inquiry into the potential utilization of quantum statistical estimation techniques to augment the theoretical convergence speeds of tabular RL algorithms within the context of infinite horizon learning settings.
@@
A diverse array of quantum statistical estimation primitives has emerged, showcasing significant enhancements in convergence speeds and gaining traction within Quantum Machine Learning (QML) frameworks (brassard2002quantum; harrow2009quantum; gilyen2019quantum). Notably, (hamoudi2021quantum) introduces a quantum mean estimation algorithm that yields a quadratic acceleration in sample complexity when contrasted with classical mean estimation techniques. In this work, we emphasize the adoption of this particular mean estimation method as a pivotal component of our methodology. It strategically refines the signals garnered by the RL agent during its interaction with the enigmatic quantum-classical hybrid MDP environment.
@@
It is pertinent to highlight a fundamental element in the analysis of conventional Reinforcement Learning (RL), which involves the utilization of martingale convergence theorems. These theorems play a crucial role in delineating the inherent stochastic process governing the evolution of states within the MDP. Conversely, this essential aspect lacks parallelism within the quantum setting, where comparable martingale convergence results remain absent. To address this disparity, we introduce an innovative approach to regret bound analysis for quantum RL. Remarkably, our methodology circumvents the reliance on martingale concentration bounds, a staple in classical RL analysis, thus navigating uncharted territories in the quantum realm.
@@
    We introduce Quantum-UCRL (Q-UCRL), a model-based, infinite horizon, optimism-driven quantum Reinforcement Learning (QRL) algorithm. Drawing inspiration from classical predecessors like the UCRL and UC-CURL algorithms (auer2008near; agarwal2021concave; agarwal2023reinforcement), Q-UCRL is designed to integrate an agent’s optimistic policy acquisition and an adept quantum mean estimator, and is the first quantum algorithm for infinite horizon average reward RL with provable guarantees.
@@
    Employing meticulous theoretical analysis, we show that the Q-UCRL algorithm achieves an exponential improvement in the regret analysis. Specifically, it attains a regret bound of $\tilde{\mathcal{O}}(1)$, breaking through the theoretical classical lower bound of $\Omega(\sqrt{T})$ across $T$ online rounds. The analysis is based on the novel quantum Bellman error based analysis (introduced in classical RL in (agarwal2021concave) for optimism based algorithm), where the difference between the performance of a policy on two different MDPs is bounded by long-term averaged Bellman error and the quantum mean estimation is used for improved Bellman error. Further, the analysis avoids dependence on martingale concentration bounds and incorporates the phenomenon of state collapse following measurements, a crucial aspect in quantum computing analysis.
  3. 3.
    We design a momentum-based estimator in equations (21)-(25) that fuses each epoch’s fresh quantum mean estimate with all past estimates using counts-based weights. The resulting algorithm preserves information that would otherwise be lost to quantum collapse, enabling accuracy to accumulate epoch-by-epoch and marking the first reuse-of-information technique fully compatible with quantum-measurement constraints. This method is essential for integrating quantum speedups into model-based RL frameworks and represents a core technical novelty of our work.
@@
Related Work and Preliminaries
Infinite Horizon Reinforcement Learning: Regret analysis of infinite horizon RL in the classical setting has been widely studied with average reward objective in both model-based settings and model-free settings (wei2020model; bai2024regret; hong2025reinforcement; ganesh2025order; ganesh2025sharper). In model-based methods, a prominent principle that underpins algorithms tailored for this scenario is the concept of “optimism in the face of uncertainty” (OFU). In this approach, the RL agent nurtures optimistic estimations of value functions and, during online iterations, selects policies aligned with the highest value estimates (fruit2018efficient; auer2008near; agarwal2021concave). Additionally, it’s noteworthy to acknowledge that several methodologies are rooted in the realm of posterior sampling, where the RL agent samples an MDP from a Bayesian Distribution and subsequently enacts the optimal policy (osband2013more; agrawal2017optimistic; agarwal2022multi; agarwal2023reinforcement). In our study, we follow the model-based approach and embrace the OFU-based algorithmic framework introduced in (agarwal2021concave), and we extend its scope to an augmented landscape where the RL agent gains access to supplementary quantum information. Furthermore, we render a mathematical characterization of regret, revealing a $\tilde{\mathcal{O}}(1)$ bound, which in turn underscores the merits of astutely processing the quantum signals within our framework.
Quantum Mean Estimation: The realm of mean estimation revolves around the identification of the average value of samples stemming from an unspecified distribution. Of paramount importance is the revelation that quantum mean estimators yield a quadratic enhancement when juxtaposed against their classical counterparts (montanaro2015quantum; hamoudi2021quantum). The key reason for this improvement is based on the quantum amplitude amplification, which allows for suppressing certain quantum states w.r.t. the states that are desired to be extracted (brassard2002quantum). In Appendix B, we present a discussion around Quantum Amplitude Estimation and its applications in QML.
@@
Definition 1 (Random Variable, Definition 2.2 of (cornelissen2022near)).
@@
The unitiary $\mathcal{U}_{\mathbb{P}}$ provides the ability to query samples of the random variable in superpositioned states, which is the essence for speedups in quantum algorithms. To perform quantum mean estimation for values of random variables (cornelissen2022near; hamoudi2021quantum; montanaro2015quantum), an additional quantum evaluation oracle would be needed.
@@
Unlike a classical sample, which discloses only one successor state, the quantum oracle stores the entire distribution in coherent superposition, offering far greater information richness per query. We note that this above form of quantum oracle, where the agent performs a quantum experiment to query the oracle and obtain a superpositioned quantum sample, is widely adopted in theoretical quantum learning literature, including optimizations (sidford2023quantum), searching algorithms (grover1996fast), and episodic RL (zhongprovably). These quantum experiment frameworks are key mechanisms that enable quantum speedups in sample complexity and regret analysis. We now present a key quantum mean estimation result that will be carefully employed in our algorithmic framework to utilize the quantum superpositioned states collected by the RL agent. One of the crucial aspects of Lemma 1 is that quantum mean estimation converges at the rate $\mathcal{O}(\frac{1}{n})$ as opposed to classical benchmark convergence rate $\mathcal{O}(\frac{1}{\sqrt{n}})$ for $n$ number of samples, therefore estimation efficiency quadratically.
Lemma 1 (Quantum multivariate bounded estimator, Theorem 3.3 of (cornelissen2022near)).
Let $X$ be a d-dimensional bounded random variable such that $||X||_{2}\leq 1$. Given three reals $L_{2}\in(0,1]$, $\delta\in(0,1)$ and $n\geq 1$ such that $\mathbb{E}[||X||_{2}]\leq L_{2}$, the multivariate bounded estimator $\texttt{QBounded}_{d}(X,L_{2},n,\delta)$ obtained by Algorithm 1 performs $f(n,d)=\mathcal{O}\Big{(}n\log^{1/2}(n\sqrt{d})\Big{)}$ quantum experiments and outputs a mean estimate $\hat{\mu}$ of $\mu=\mathbb{E}[X]$ such that,
@@
Algorithm 1 $\texttt{QBounded}_{d}(X,L_{2},n,\delta)$ (Algorithm 1 in (cornelissen2022near))
@@
8:     Compute the state $|\psi\rangle:=\widetilde{P}_{X,L_{2},m,\alpha,\epsilon}|G\rangle$ in $H_{G}\otimes H_{aux}$, where $\widetilde{P}_{X,L_{2},m,\alpha,\epsilon}$ is the directional mean oracle constructed by the quantum evaluation oracle in Proposition 3.2 (cornelissen2022near) with $\epsilon=1/25$.
@@
Quantum Reinforcement Learning: Within the realm of QRL, a prominent strand of previous research showcases exponential enhancements in regret through amplitude amplification-based methodologies applied to the Quantum Multi-Armed Bandits (Q-MAB) problem (wang2021quantum; casale2020quantum; wan2023quantum; wu2023quantum). Nevertheless, the theoretical underpinnings of the aforementioned Q-MAB approaches do not seamlessly extend to the QRL context since there is no state evolution in bandits.
A recent surge of research interest has been directed towards Quantum Reinforcement Learning (QRL) (jerbi2021quantum; dunjko2017advances; paparo2014quantum). It is noteworthy to emphasize that the aforementioned studies do not provide an exhaustive mathematical characterization of regret. (zhongprovably; ganguly2023quantum) demonstrated that QRL can provide logrithmic regret in the episodic MDP settings. Ours is the first study of infinite horizon MDP with average reward objective.
Problem Formulation
@@
Quantum Computing Basics: In this section, we give a brief introduction to the concepts that are most related to our work based on (nielsen2010quantum) and (wu2023quantum). Consider an $m$ dimensional Hilbert space $\mathbb{C}^{m}$, a quantum state $|x\rangle=(x_{1},...,x_{m})^{T}$ can be seen as a vector inside the Hilbert space with $\sum_{i}|x_{i}|^{2}=1$. Furthermore, for a finite set with $m$ elements $\xi=\{\xi_{1},...,\xi_{m}\}$, we can assign each $\xi_{i}\in\xi$ with a member of an orthonormal basis of the Hilbert space $\mathbb{C}^{m}$ by
@@
Quantum transition oracle: We utilize the concepts and notations of quantum computing in (wang2021quantum; wiedemann2022quantum; ganguly2023quantum; jerbi2022quantum) to construct the quantum sampling oracle of RL environments and capture agent’s interaction with the unknown MDP environment. We now formulate the equivalent quantum-accessible RL environments for our classical MDP $\mathcal{M}$. For an agent at step $t$ in state $s_{t}$ and with action $a_{t}$, we construct the quantum sampling oracles for variables of the next state $P(\cdot|s_{t},a_{t})$. Specifically, suppose there are two Hilbert spaces $\mathcal{\bar{S}}=\mathbb{C}^{|\mathcal{S}|}$ and $\mathcal{\bar{A}}=\mathbb{C}^{|\mathcal{A}|}$ containing the superpositions of the classical states and actions. We represent the computational bases for $\mathcal{\bar{S}}$ and $\mathcal{\bar{A}}$ as $\{|s\rangle\}_{s\in\mathcal{S}}$ and $\{|a\rangle\}_{a\in\mathcal{A}}$. We assume that we can implement the quantum sampling oracle $\mathcal{U}_{P}$ of the MDP’s transitions as follows:
@@
$\mathcal{U}_{P}:|s_{t}\rangle\otimes|a_{t}\rangle\otimes|0\rangle\rightarrow|\bf{\psi}\rangle^{t}$
Where $|\bf{\psi}\rangle^{t}=|s_{t}\rangle\otimes|a_{t}\rangle\otimes\sum_{s^{\prime}\in\mathcal{S}}\sqrt{P(s^{\prime}|s_{t},a_{t})}|s^{\prime}\rangle$.
@@
As described in Fig 1, at step $t$, the agent plays an action $a_{t}$ based on current state $s_{t}$ according to its policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$. Consequently, the unknown MDP Q-environment shares the next state information and reward i.e., $\{s_{t+1},r_{t}\}$. Additionally, the quantum transition oracle first encodes $s_{t}$ and $a_{t}$ into the corresponding basis of the quantum Hilbert spaces, producing basis $|s_{t}\rangle$ and $|a_{t}\rangle$, then it encodes the superpositioned quantum state $|\bf{\psi}\rangle^{t}$ from the quantum transition oracle $\mathcal{U}_{P}$. $|\bf{\psi}\rangle^{t}$ would be used for the Quantum Mean Estimator, QBounded, which in turn improves transition probability estimation and leads to exponentially better regret. We note that one copy of $|\bf{\psi}\rangle^{t}$ is obtained at each $(s_{t},a_{t})$. Given that a measurement collapses the state, we can only perform one measurement on it. So, the quantum mean estimator will only perform measurements at epoch ends, as will be described in the algorithm.
RL regret formulation: Consider the RL agent following some policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$ on the unknown MDP Q-environment $\mathcal{M}$. Then, as a result of playing policy $\pi$, the agent’s long term average reward is defined next according to (agarwal2021concave).
@@
Then, according to (puterman2014markov), $\Gamma_{\pi}^{P}$ can be represented as the following:
@@
wherein, $V_{\pi}^{P}(\cdot;\cdot)$ is the discounted cumulative average reward by following policy $\pi$, $\rho_{\pi}^{P}$ is the steady-state occupancy measure (i.e., $\rho_{\pi}^{P}(s,a)\in[0,1],~{}\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\rho_{\pi}^{P}(s,a)=1$), $\bar{r}(s,a)$ is the steady-state long-term average reward of pair $(s,a)$. In the following, we introduce key assumptions crucial from the perspective of our algorithmic framework and its analysis. In this context, we introduce the notations $\{P_{\pi,s}^{t},T_{\pi,~{}s\rightarrow s^{\prime}}\}$ to denote the $t$ time step probability distribution obtained by playing policy $\pi$ on the unknown MDP $\mathcal{M}$ starting from some arbitrary state $s\in\mathcal{S}$; and, the actual time steps to reach $s^{\prime}$ from $s$ by playing policy $\pi$ respectively. Next, we state an assumption pertaining to ergodicity of MDP $\mathcal{M}$.
@@
We emphasize that Assumption 2 is a commonly used assumption in RL algorithm development owing to the fact that in most setups rewards are constructed according to the underlying problem and is known beforehand (azar2017minimax; agarwal2019reinforcement). Next, we define the cumulative regret accumulated by the agent in expectation across $T$ time steps as follows:
@@
However, in the absence of knowledge pertaining to true $P$, we propose to solve an optimistic version of Eq. (11) that will update RL agent’s policy in epochs. Specifically, at the start of an epoch $e$, the following optimization problem will be solved:
@@
Observe that Eq. (15) increases the search space for policy in a carefully computed neighborhood of certain transition probability estimates $\hat{P}$ via Eq. (19) at the start of epoch $e$. Also in Eq. (19), recall that $S=|\mathcal{S}|$ is the number of states, $A=|\mathcal{A}|$ is the number of actions, $N_{e}(s,a)$ is the number of visitations of state-action pair $(s,a)$ upto the end of epoch $e$. Note that, solving Eq. (15) is equivalent to obtaining an optimistic policy as it extends the search space of original problem Eq. (11). Furthermore, the computation of $\hat{P}_{e}$ is key to our framework design. Since performing quantum mean estimation to quantum samples would require measuring and results in quantum samples collapsing, each quantum sample can only be used once. We emphasize that the quantum mean estimator will only run at the end of the epoch, performing measurements to update the transition probability for the next epoch.
Details on Quantum Mean Estimator: At the end of the epoch, we use all $(s_{i},a_{i})$ observed in the samples, and perform quantum mean estimation on all the samples $|\bf{\psi}\rangle^{t}$ obtained for that $(s_{i},a_{i})$. More precisely, $\tilde{P}_{e}(\cdot|s_{i},a_{i})$ is the quantum mean estimation for the random variable $P_{e}(\cdot|s_{i},a_{i})$ performed by QBounded (Algorithm 1) using the quantum samples $\{|{\psi}\rangle\}^{e}_{i}$ observed in the previous epoch with the total number being $\nu_{e-1}(s_{i},a_{i})$:
${\{|{\psi}\rangle\}^{e}_{i}=\{|{\psi}\rangle^{t}\,|\;t_{e-1}^{start}\leq t\leq t_{e}^{start}-1,(s_{t},a_{t})=(s_{i},a_{i})\}}$ (20)
@@
for $c\in\mathbb{R}$ satisfying $cn\log^{1/2}(n\sqrt{S})\geq f(n,S)\;\forall n$ where $f(n,S)$ is defined in Lemma 1. Such choice of $c$ would allow $\nu_{e}(s,a)$ samples to be used for analysis of our proposed algorithm. Note that the choice of $c$ is achievable because $cn\log^{1/2}(n\sqrt{S})$ and $f(n,S)$ are the same order. Since $\mathbb{E}[||P_{e}(\cdot|s_{i},a_{i})||_{2}]\leq 1$, we set $L_{2}=1.$ Different from the works in classical approaches based on (jaksch2010near), our algorithm utilizes each quantum sample only once for estimating all $\hat{P}_{e}$ throughout the entire process. This strategy fits with our quantum framework since quantum samples would collapse once being measured. In Appendix A, we show that Eq. (15) is a convex optimization problem, therefore standard optimization solvers can be employed to solve it. Finally, the agent’s policy at the start of epoch $e$ is given by:
@@
Next, we formally present Q-UCRL in Algorithm 2.
@@
11:     Observe $s_{t+1},r(s_{t},a_{t})$ and quantum samples $|\bf{\psi}\rangle^{t}$.